# Bridging the Long-Term Gap: A Memory-Active Policy for Multi-Session Task-Oriented Dialogue

### v1

@@ Abstract
Existing Task-Oriented Dialogue (TOD) systems primarily focus on single-session dialogues, limiting their effectiveness in long-term memory augmentation. To address this challenge, we introduce a MS-TOD [^1] Code and dataset will be released upon paper acceptance. dataset, the first multi-session TOD dataset designed to retain long-term memory across sessions, enabling fewer turns and more efficient task completion. This defines a new benchmark task for evaluating long-term memory in multi-session TOD. Based on this new dataset, we propose a Memory-Active Policy (MAP) that improves multi-session dialogue efficiency through a two-stage approach. 1) Memory-Guided Dialogue Planning retrieves intent-aligned history, identifies key QA units via a memory judger, refines them by removing redundant questions, and generates responses based on the reconstructed memory. 2) Proactive Response Strategy detects and correct errors or omissions, ensuring efficient and accurate task completion. We evaluate MAP on MS-TOD dataset, focusing on response quality and effectiveness of the proactive strategy. Experiments on MS-TOD demonstrate that MAP significantly improves task success and turn efficiency in multi-session scenarios, while maintaining competitive performance on conventional single-session tasks.
Bridging the Long-Term Gap: A Memory-Active Policy for
Multi-Session Task-Oriented Dialogue
Yiming Du1thanks: These authors contributes equally., Bingbing Wang211footnotemark: 1, Yang He3, Bin Liang1, Baojun Wang4, Zhongyang Li4, Lin Gui5, Jeff Z. Pan6, Ruifeng Xu2, Kam-Fai Wong1 1The Chinese University of Hong Kong 2Harbin Institute of Technology, Shenzhen 3 HKUST 4Huawei Noah’s Ark Lab 5King’s College London 6 The University of Edinburgh {ydu, kfwong}@se.cuhk.edu.hk, {bingbing.wang}@stu.hit.edu.cn
1 Introduction
Task-oriented dialogue (TOD) systems Wang et al. (2021); He et al. (2022); Bang et al. (2023); Swamy et al. (2023a) have traditionally focused on single-session scenarios, overlooking the fact that real world interactions often span multiple sessions over extended periods. While LLMs have been introduced to improve TOD Xu et al. (2024a, b); Chung et al. (2023); Heck et al. (2023a), most efforts remain confined to single-session settings and overlook long-term memory augmentation across multi-session interactions (Du et al., 2025). Moreover, existing TOD datasets Stacey et al. (2024); Liu et al. (2024a); Budzianowski et al. (2018); Rastogi et al. (2020) are limited to single-session dialogues, leaving a gap in benchmarks for evaluating long-term memory retention across sessions.
@@ 1 Introduction
Figure 1: Task-oriented dialogue, without (left) vs. with (right) multi-session memory; the latter demands more turns of conversation.
As shown in Figure 1, single-session systems require users to restate details (e.g., flight times, seat preferences) in every session, leading to inefficiency and frustration. In contrast, multi-session memory enables seamless retrieval of prior context, supporting fewer turns and a more personalized experience. In contrast, multi-session memory allows systems to retrieve prior user-specific information—such as preferences or booking history—thereby reducing redundancy and supporting more efficient, goal-oriented interactions.
To bridge this gap, we introduce the Multi-session Task-oriented Dialogue Dataset (MS-TOD), comprising 132 simulated speaker partitions, each spanning over 20 sessions with diverse task goals derived from SGD Rastogi et al. (2020). MS-TOD supports comprehensive evaluation of TOD systems to retrieve long-term context, maintain consistent task slots, and adapt responses across sessions. This also defines a novel task setting for evaluating long-term memory integration in multi-session TOD, where systems must leverage cross-session knowledge to support coherent and efficient task completion. While existing approaches in open-domain multi-session conversations focus on retrieving dialogue history or summaries Lu et al. (2023); Zhong et al. (2024); Joko et al. (2024); Li et al. (2024a); Du et al. (2024), multi-session TOD system face additional demands: they must recall critical slot-value pairs, track evolving user intents, and proactively resolve missing or outdated information while minimizing redundant queries.
Therefore, we propose the Memory-Active Policy (MAP) to incorporate long-term memory in the multi-session TOD task. MAP consists of two core phases: (1) Memory-Guided Dialogue Planning, where an LLM generates an intent hypothesis retrieves relevant memory entries to support cross-session goal tracking. Furthermore, a memory judger identifies key QA units and refines them by removing redundant questions, distilling slot-level content for precise, context-aware response generation. (2) Proactive Response Strategy, which iteratively detects missing or mismatched slots by comparing predicted responses with task goals, actively engaging users to resolve incomplete slots, thereby reducing redundancy and ensuring smooth, goal-oriented interactions. Experimental results on MS-TOD demonstrate that MAP effectively improves dialogue coherence, response quality, task success rate, and dialogue efficiency in multi-session TOD. The main contributions include:
@@
  • We propose MAP, a two-stage framework that distills and leverages cross-session memory for efficient, minimal-turn task completion.
  • We demonstrate that MAP consistently outperforms strong baselines across multiple metrics, confirming the effectiveness of its memory activation policy.
@@
Full Conversation Context 4.76 0.61
[…]
@@
2 Preliminary Experiments
To investigate the effectiveness of different strategies for handling dialogue history in multi-session task-oriented response generation, We conduct a preliminary study comparing direct prompting Swamy et al. (2023b); Xu et al. (2024a) with retrieval-augmented generation (RAG) Huang et al. (2024); Lu et al. (2023) in multi-session TOD.
Because standard TOD datasets lack multi-session dependencies, we construct a test set specifically for confirmation-type response generation (details in Section 3). Our pipeline includes (1) Retrieval. We explore three strategies for retrieving relevant historical dialogues: sparse retrieval (BM25 Robertson and Zaragoza (2009)), dense retrieval (text-embedding-small-3222OpenAI. text-embedding-3-small. 2025. OpenAI, https://platform.openai.com/docs/guides/embeddings.), and a hybrid approach that combines both to leverage their complementary strengths. (2) Response Generation. GPT-4o-mini then generates confirmation-type responses by incorporating the retrieved dialogue history and task goal information.
As shown in Table 1, RAG consistently outperforms direct prompting. For instance, dense retrieval achieves 0.67 slot accuracy and a 7.01 GPT-4 score, surpassing full-context prompting (0.61 and 4.76, respectively). Hybrid retrieval further improves slot accuracy to 0.68 and the GPT-4 score to 7.04, demonstrating the value in combining sparse and dense strategies. Oracle retrieval (using ground-truth context) reaches 0.88 and 8.51, underscoring the need for more accurate retrieval strategies in multi-session TOD.
@@ 2 Preliminary Experiments
3 Dataset
To systematically evaluate TOD systems requiring multi-session long-term memory integration, we develop the MS-TOD dataset, derived from the Schema-Guided Dialogue (SGD) dataset Rastogi et al. (2020). Compared with existing TOD and open-domain datasets, MS-TOD uniquely supports multi-session memory retrieval, slot tracking, and intent continuity (see Appendix Table 9 for dataset comparison). MS-TOD comprises two subsets: a training subset for training the memory judger (Section 4.2) and an evaluation subset designed to assess multi-session memory activation and TOD response generation shown in Figure 2.
3.1 Data Generation
Multi-Session Dialogue Construction. Because existing TOD corpora typically feature single-session interactions lacking structured multi-session dependencies, we create three dialogue sessions for each task in the SGD dataset. Compared with single-session dialogues, this design more closely simulates how users revisit and refine the same task at different times and in different contexts. We chose three sessions—rather than a higher number—to strike a balance between capturing realistic user behavior and avoiding repetitive dialogue data, particularly given that SGD tasks involve fewer than ten task slots. As a result, three sessions offer sufficient coverage of task variations without overpopulating the dataset. More details can be found in Appendix A.1.
Confirmation-Type Response Annotation. In the final session of each task, we introduce confirmation-type annotations to mark utterances indicating the completion of long-term or recurring tasks. These annotations serve two primary functions: (1) Guiding Memory Activation: Highlighting key dialogue points to trigger long-term memory activation, summaries, or confirmations; and (2) Supporting System Evaluation: They enable evaluation of the system’s ability to recognize and record cross-session information or long-term goals during dialogue strategy assessment.
@@ 3.1 Data Generation
Figure 3: Overflow of our MAP framework, which comprises Memory-Guided Dialogue Planning and Proactive Response Strategy.
3.2 Memory Bank Construction
Since multi-session interactions occur at the individual level, we group sessions into Individual Memory Banks (Figure 2), each storing an individual’s historical dialogues to maintain continuity and enable adaptive responses. Each bank contains over 20 sessions spanning more than six distinct user intentions (Table 2), with one evaluation session per intent to assess confirmation-type generation. Task goals are included to support memory activation and task handling. We employ a GPT-4-based generator to extract intent descriptions and construct task-specific QA pairs (Appendix A.2), enabling efficient and intention-aware memory retrieval across domains. Additional dataset details appear in Appendix B.1 and B.2.
MS-TOD is derived from the Schema-Guided Dialogue (SGD) dataset, selected for its broad domain coverage, schema-driven design, and support for multi-domain interactions. Our memory bank structure (20 sessions, 6+ intents) reflects typical slot-intent patterns in SGD and aligns with prior multi-session benchmarks (Appendix B.5).
Human Validation. To ensure the quality and coherence of the constructed dataset, we conducted a multi-stage manual validation process to verify intent accuracy, slot-value correctness, dialogue coherence, and confirmation-response alignment. Full validation procedures are described in Appendix B.3.
4 Memory-Active Policy
To address the need for long-term memory and multi-session context in TOD, the Memory-Active Policy (MAP) combines memory-driven dialogue planning with a proactive policy strategy as shown in Figure 3.
4.1 Task Definition
The objective of this task is to generate a natural language response $r$ based on the provided dialogue context $c$ and individual memory bank $M$. The dialogue context $c$ represents the ongoing interaction, comprising chronologically ordered user utterances $u_{j}$ and system responses $r_{j}$. The individual memory bank $M$ consists of memory representations from multiple dialogue sessions, where each session provides an intent description $k_{i}$ and a corresponding set of QA pairs $v_{i}$. Formally, we define: $M=\{(k_{i},v_{i})\}_{i=1}^{N},v_{i}=\{(q_{ij},a_{ij})\}_{j=1}^{n}.$ Here, $k_{i}$ represents the session’s high-level intent, while $v_{i}$ stores detailed task-related information. This structured design facilitates efficient retrieval and utilization of long-term user-specific memory.
The response $r$ is generated by a large language model $LLM$ that integrates $c$ and $M$, ensuring semantic coherence, memory relevance, and task slot accuracy. Formally, the optimal response is obtained by maximizing the conditional probability distribution:
$r^{*}=\arg\max_{r\in\mathcal{R}}P(r\mid c,M),$ (1)
where $\mathcal{R}$ denotes the set of all possible responses. This approach emphasizes understanding the dialogue context and leveraging individual memory to produce coherent and relevant responses.
4.2 Memory-Guided Dialogue Planning
Memory-Guided Dialogue Planning consists of two key steps: (1) Intent Capture and Memory Activation, where the system identifies and retrieves relevant memories aligned with the user’s intentions; (2) Memory Judgement and Refinement, which detects missing task slots, and re-ranks relevant memories to ensure optimal information recall for response generation.
Intent Capture and Memory Activation.
Given the dialogue context $c_{i}$, we use LLM (GPT4o-mini) to generate a high-level intent description $k_{i}$, which summarizes the user’s objective in the current session. The intent description $k_{i}$ is then used to retrieve relevant memory units from the long-term memory $M$, represented as $M=\{(k_{j},v_{j})\}_{j=1}^{m}$, where $k_{j}$ is an intent-related key and $v_{j}$ is the corresponding structured information, such as paired questions and answers. Using an embedding model, $k_{i}$ is mapped to a dense representation and compared with $k_{j}$ to activate the most relevant memory units $v_{i}$. These activated memory units $v_{i}$, containing structured information such as task-related questions and answers, are then used to guide subsequent dialogue processing.
Memory Judgement and Refinement.
To guide accurate memory selection, we first identify missing task information using a Chain-of-Thought (CoT) prompting strategy Wei et al. (2022), which generates a hypothesized missing slot query $q_{\text{miss}}$ based on the current dialogue context $c_{i}$ This missing query serves as a proxy for what information is needed from memory to complete the task. We then assess the relevance of each QA pair in the retrieved memory using a memory judger instantiated by LLaMA 3.1-8B (Meta AI, 2024). Given $c_{i}$, $q_{\text{miss}}$ and $(q_{j,u},a_{j,u})$, the judger estimates:
$s_{j,u}=P(y=1|(c_{i},q_{\text{miss}},(q_{j,u},a_{j,u})))$ (2)
where $y=1$ indicates that the QA pair contributes to the task goal, while $y=0$ indicates irrelevance. The model is trained with cross-entropy loss:
$\mathcal{L}=-\sum y\log s_{j,u}+(1-y)\log(1-s_{j,u})$ (3)
To balance retrieval relevance and judger evaluation, we re-rank all QA pairs using a weighted combination of initial retrieval score $s_{\text{f},ju}$ and judger score $s_{ju}$:
$s_{\text{f},ju}=\alpha\cdot s_{\text{prev},ju}+(1-\alpha)\cdot s_{ju},$ (4)
@@ 4.2 Memory-Guided Dialogue Planning
The top 5 QA pairs $v_{selected}$ with the highest $s_{\text{final},ju}$ scores are selected for refinement.
Memory Refinement. In the refinement stage, we clean and restructure the selected QA pairs to remove noise and improve response grounding. Specifically, we discard auxiliary questions $q_{j,u}$) and retain only the core answers $A_{\text{core}}=\{a_{j,u}\}$. These core answers are concatenated into the dialogue context to form a memory-enhanced prompt for the generation module. This step ensures that only high-relevance, low-noise content is passed forward for response planning.
4.3 Proactive Response Strategy
The response generation phase synthesizes the dialogue context $c$ and pruned memory $A_{\text{core}}$ (from memory reconstruction) into a confirmation response $r$. Using $\text{LLM}_{\text{Reader}}$, the system evaluates if integrated memory supports task completion:
$r=\text{LLM}_{\text{Reader}}(c,A_{\text{core}})$ (5)
where $r$ serves dual purposes: (1) providing task guidance and (2) explicitly verifying memory relevance to user goals (see Appendix A.3 for details).
To provide more comprehensive responses to user queries, we propose a proactive dialogue policy. Based on the generated response $r$, we identify missing or incorrect slots within the dialogue. This results in a set of slots, denoted as $L=\{l_{1},l_{2},...,l_{n}\}$, where each $l_{i}$ represents a missing or erroneous slot. We design an agent to simulate the user, explicitly informing it of the slot set $L$. The user agent then interacts with our dialogue model in an interactive conversation to address the identified slots.
At each dialogue turn, a supervisor (played by an LLM) evaluates whether the conversation accurately fulfills the slot information requirements. If a slot $s_{i}$ is successfully resolved during the interaction, it is removed from $L$. Mathematically, the update to the slot set is expressed as:
$S\leftarrow L\ \{l_{i}\}$ (6)
The interaction continues for multiple turns until the slot set becomes empty, $L=\emptyset$, ensuring all missing or erroneous slots are resolved.
5 Experiments
5.1 Experimental Setups

### v2

# MemGuide: Intent-Driven Memory Selection for Goal-Oriented Multi-Session LLM Agents

@@ Abstract
Modern task-oriented dialogue (TOD) systems increasingly rely on large language model (LLM) agents, leveraging Retrieval-Augmented Generation (RAG) and long-context capabilities for long-term memory utilization. However, these methods are primarily based on semantic similarity, overlooking task intent and reducing task coherence in multi-session dialogues. To address this challenge, we introduce MemGuide, a two-stage framework for intent-driven memory selection. (1) Intent‑Aligned Retrieval matches the current dialogue context with stored intent descriptions in the memory bank, retrieving QA‑formatted memory units that share the same goal. (2) Missing‑Slot Guided Filtering employs a chain‑of‑thought slot reasoner to enumerate unfilled slots, then uses a fine‑tuned LLaMA‑8B filter to re‑rank the retrieved units by marginal slot‑completion gain. The resulting memory units inform a proactive strategy that minimizes conversational turns by directly addressing information gaps. Based on this framework, we introduce the MS-TOD [^1] Code and dataset will be released upon paper acceptance., the first multi-session TOD benchmark comprising 132 diverse personas, 956 task goals, and annotated intent-aligned memory targets, supporting efficient multi‑session task completion. Evaluations on MS-TOD show that MemGuide raises the task success rate by 11% (88%→99%) and reduces dialogue length by 2.84 turns in multi-session settings, and maintains parity with single‑session benchmarks.
@@
Figure 1: Task-oriented dialogue, without (left) vs. with (right) multi-session memory; the former demands more turns of conversation.
Introduction
Modern task-oriented dialogue (TOD) systems increasingly integrate large language models (LLMs) to enhance generalization, context understanding, and response generation (Nguyen et al. 2025; Xu et al. 2024a, b; Chung et al. 2023; Hudeček and Dusek 2023). To utilize historical dialogue states across turns, two dominant strategies have emerged: retrieval-augmented generation (RAG), which supplements the LLM with relevant task descriptions or prior dialogue states (Xu et al. 2024a, b), and the use of long-context models, which encode the entire history directly as input (Nguyen et al. 2025). However, they are both limited to surface-level semantic similarity, often ignoring the task-specific intent and slot-level continuity crucial for coherent multi-session TOD.
While memory-augmented methods (Lu et al. 2023) have emerged, they are often evaluated on single-session benchmarks that lack support for long-term goal tracking and multi-session memory supervision. Unlike open-domain settings that focus on free-form recall (Zhong et al. 2024), task-oriented dialogue demands structured slot tracking, evolving intent management, and consistent state maintenance. Crucially, real-world users frequently interact with assistants across multiple sessions to accomplish complex goals, yet most existing TOD models and datasets (Budzianowski et al. 2018; Rastogi et al. 2020; Stacey et al. 2024; Liu et al. 2024a) are confined to single-session settings, highlighting a fundamental gap between current TOD systems and the demands of persistent, goal-driven dialogue settings. As shown in Figure 1, traditional single-session TOD systems require users to restate details (e.g., flight times, seat preferences) in every session, leading to inefficiency and frustration.
To address this, we propose MemGuide, a novel framework that utilizes long-term memory in the multi-session TOD task. Unlike prior methods that rely solely on semantic similarity for memory selection, MemGuide incorporates task intent and slot-level guidance to enhance memory relevance and response quality. It consists of two core phases: (1) Intent-Aligned Retrieval, where an LLM generates an intent hypothesis based on the current dialogue state and retrieves memory units aligned with the predicted goal, ensuring retrieved history is both semantically similar and goal-consistent to support cross-session task tracking robustly. (2) Missing-Slot Guided Filtering, which first employs a chain-of-thought (CoT) slot reasoner to detect missing slot values, followed by a filtering module that removes irrelevant or redundant Question Answering (QA) memory units to distill slot-level content for response generation. These two phases enable MemGuide to convert long-term user history into actionable context, supporting minimal-turn, task-consistent dialogue generation across sessions.
To enable systematic evaluation, we construct the Multi-Session Task-oriented Dialogue (MS-TOD), a new benchmark comprising 132 simulated speakers, each engaging in over 20 sessions covering diverse task goals derived from Schema-Guided Dialogue (SGD) (Rastogi et al. 2020). MS-TOD supports evaluation of slot continuity and long-term memory retrieval across sessions. Unlike open-domain benchmarks focused on retrieving dialogue summaries (Zhong et al. 2024; Li et al. 2024a; Du et al. 2024), multi-session TOD introduces additional challenges. Systems must recall key slot-value pairs, track evolving user intents, and proactively resolve missing or outdated information, while minimizing redundant interactions. To support automatic evaluation, we introduce a proactive response generation module to simulate user engagement and evaluate system performance in resolving missing information. Experimental results demonstrate that MemGuide significantly improves dialogue coherence, response quality, task success rate, and overall efficiency in multi-session TOD. The main contributions include:
  • We propose MemGuide, a two-stage framework that distills and leverages cross-session memory for efficient, minimal-turn task completion.
@@ Introduction
  • We demonstrate that MemGuide consistently outperforms strong baselines across multiple metrics, demonstrating the effectiveness of intent-aware memory retrieval and slot-guided filtering.
Related Work
Task-Oriented Dialogue Dataset
TOD datasets are typically constructed via either Machine-to-Machine (M2M) (Shah et al. 2018; Rastogi et al. 2020) or Wizard-of-Oz (WOz) setups (Wen et al. 2017; Budzianowski et al. 2018). M2M datasets (e.g., SGD, STAR) provide schema-driven task flows, while WOz-based datasets (e.g., MultiWOZ, FRAMES) offer more natural but annotation-heavy dialogues. Recent efforts aim to improve realism and domain diversity (Hu et al. 2023; Dai et al. 2022; Xu et al. 2024b; Li et al. 2024b), yet existing benchmarks primarily assume single-session tasks. There remains a notable gap in datasets designed for multi-session TOD, where tracking long-range goals and user intents is essential.
Task-Oriented Dialogue Systems
Traditional TOD systems adopt modular pipelines for NLU, DST, and response generation (Wu et al. 2019a; Peng et al. 2018), later unified into end-to-end models trained on annotated dialogues (Wen et al. 2017; Wang et al. 2020). With the rise of LLMs, recent work explores their use in zero-shot and fine-tuned TOD (Madotto et al. 2021; Bang, Lee, and Koo 2023), often achieving strong results on intent recognition and slot filling. In parallel, long-term memory (LTM) methods such as ChitChat (Li et al. 2024c), MemoryBank (Zhong et al. 2024), and LoCoMo (Maharana et al. 2024) support extended context retention through summarization or heuristic filtering, but lack structured memory aligned with task goals. Most assume single-session dialogues and overlook challenges in maintaining multi-session goal continuity. This work addresses these gaps by introducing MemGuide for long-range, goal-aware tracking.
@@
Full Context 4.76 0.61
Retrieval-based Methods
@@
Preliminary Experiments
To motivate our framework, we first examine the limitations of direct prompting in multi-session TOD and explore the potential of retrieval-based strategies. Since existing TOD datasets lack long-term dependencies, we construct an evaluation set focused on confirmation-type response generation. This evaluation set adopts the same input–output formulation, evaluation criteria, and sample set as the MS-TOD benchmark to ensure consistency and comparability.
To clarify the evaluation setting, we formulate a confirmation-style response generation task. Formally, given the current user utterance $u_{t}$, the dialogue context $c=\{u_{1},r_{1},\dots,u_{t-1},r_{t-1}\}$ from the current session, and history dialogue sessions $H$ in the corresponding memory bank (retrieved or concatenated), the model generates a confirmation-style response $r_{t}$. The task goal $g$ is held out for reference during evaluation. This setting mimics a common scenario in TOD where the system determines whether the dialogue contains sufficient information to proceed with task execution. For all settings, we use GPT-4o-mini as the unified generator, allowing a fair comparison across input strategies. We compare two strategies:
(1) Retrieval-based methods, including sparse (BM25 (Robertson and Zaragoza 2009)), dense (text-embedding-small-3222OpenAI. text-embedding-3-small. 2025. https://platform.openai.com/docs/guides/embeddings), and a hybrid retrieval. Each selects top-$k$ history sessions $H$ relevant to the current utterance $u_{t}$ with the model input $x=[H;c;u_{t}]$. (2) Direct prompting, where the full dialogue history is concatenated with the current user utterance $u_{t}$ without retrieval. The input is $x=[H;C;u_{t}]$. The model generates a confirmation response $r_{t}=\text{LLM}(x)$.
As Table 1 shows, retrieval-based methods consistently outperform direct prompting. For instance, dense retrieval-based method achieves 0.67 slot accuracy and 7.01 GPT-4 score, surpassing full-context prompting (0.61 and 4.76). This significant gap highlights how direct prompting struggles with context limitations and the ”lost-in-the-middle” (Liu et al. 2023), where irrelevant history overwhelms key information, motivating our development of MemGuide for advanced intent-aligned retrieval.
@@
Dataset
MS-TOD is a multi-session benchmark for evaluating long-term memory through user-specific memory banks, with over 20 sessions from a single user. This supports assessment of memory retrieval, slot tracking, and intent continuity. For evaluation, we include held-out sessions with manually annotated confirmation-type responses (Figure 2). MS-TOD is built in four stages: (1) multi-session dialogue generation; (2) confirmation-type response annotation; (3) QA-style memory bank construction; and (4) human validation.
Multi-Session Dialogue Generation
We begin by generating multi-session dialogues for each task goal sampled from the SGD dataset (Rastogi et al. 2020). For every task, we synthesize three temporally ordered sessions using GPT-4, each conditioned on the slot-filling status of the previous session. This simulates how users revisit and revise the same task across time. Specifically, Session 1 presents an incomplete task with missing slots; Session 2 introduces updates; and Session 3 concludes with final confirmation. This staged construction reflects real-world dialogue dynamics while avoiding overfitting, as most SGD tasks involve fewer than ten slots.
Confirmation-Type Response Annotation
To evaluate long-term task fulfillment in dialogue systems, we annotate the final session of each task with confirmation-type responses. Each marks the utterance confirming task completion and associated slot-value goal with manually labels (confirmation/non-confirmation). These annotations serve two purposes: (1) Supervising Memory Selection, indicating when to trigger memory retrieval; and (2) Supporting Evaluation, evaluating if the system recalls goal-relevant content and generates accurate confirmations.
@@
Figure 3: Overflow of our MemGuide framework, which comprises Intent-Aligned Retrieval and Missing-Slot Guided Filtering.
Memory Bank Construction
Since multi-session interactions are organized around individuals, we group dialogues into individual memory banks (Figure 2), where each bank stores temporally ordered sessions. During construction, we ensure that intents within the same bank are distinct and non-conflicting, enabling consistent memory usage and avoiding cross-intent interference. Each bank contains over 20 sessions spanning at least six task intents (Table 2), with one held-out evaluation session per intent for confirmation-type assessment. For each completed session, we use GPT-4 to generate an intent description along with a set of QA pairs, each capturing a slot-specific fact for retrieval. This QA-style format is motivated by prior work (Chen et al. 2023) showing that question–answer structures facilitate more accurate and efficient retrieval compared to unstructured text. To maintain causal consistency, memory access is limited to sessions prior to the current evaluation point.
Human Validation
To ensure coherence, correctness, and usability, we apply a structured multi-stage validation process involving three annotators experienced in NLP. The process includes: (1) Verifying each session preserves the intended goal and correct slot values; (2) Removing dialogues with excessive redundancy across sessions; (3) Verifying that confirmation-type utterances match the expected slot-value goals; (4) Removing sessions that fail to complete defined task goals or lack confirmation turns; and (5) Excluding episodes with unnatural repetition of similar intents. We additionally conduct inter-annotator agreement evaluation to ensure labeling consistency.
MemGuide
MemGuide is a two-stage framework for multi-session TOD, as shown in Figure 3. It first performs Intent-Aligned Retrieval, extracting the current user intent via an LLM and retrieving QA-formatted memory units from the memory bank that align with this intent using semantic similarity. Then, Missing-Slot Guided Filtering identifies unfilled task slots through a CoT reasoner and re-ranks retrieved memories with a fine-tuned LLaMA-8B filter based on their ability to fill these slots. Finally, Response Generation leverages the filtered memories to produce proactive responses, reducing conversational turns and boosting task success.
Given a dialogue context $c=\{u_{1},r_{1},\dots,u_{t-1},r_{t-1},u_{t}\}$ and a user-specific long-term memory bank $M$, the goal is to generate the next system response $r_{t}$. The memory bank $M$ is a collection of memories from past sessions, structured as $M=\{(k_{i},V_{i})\}_{i=1}^{N}$ and $t$ is the number of turns. Each entry consists of a high-level intent description $k_{i}$ (e.g., ”book a flight to San Francisco”) represents the high-level intent and a corresponding set of QA-formatted memory units $V_{i}=\{(q_{i,j},a_{i,j})\}_{j=1}^{n}$, where $q_{i,j}$ is the $j$-th question about a task slot within the $i$-th intent (e.g., ”What is the departure date?”) and $a_{i,j}$ is the corresponding answer (e.g., ”July 28, 2025”). The optimal response $r$ should coherently continue the conversation while strategically utilizing information from $M$ to progress towards task completion. It is generated by an LLM conditioned on the context and a carefully selected memory subset $M_{sel}\subseteq M$.
Intent-Aligned Retrieval
The stage aims to broadly identify past sessions relevant to the current user goal by retrieving memory entries from the long-term bank that share a consistent high-level intent, ensuring thematic alignment.
Current Intent Extraction.
Given the dialogue context $c$, we use GPT-4o-mini to generate a high-level intent description $d_{int}$, which summarizes the user’s objective in the current session. The model is prompted to generate a short, command-like phrase summarizing the user’s goal (e.g., ”find a local Italian restaurant”). This distilled phrase serves as a current intent key, $k_{cur}$, providing a canonical task representation that is standardized for retrieval.
Semantic Retrieval.
We then retrieve memory units from the bank $M$, which contain stored intent keys $\{k_{i}\}_{i=1}^{N}.$ and corresponding QA pairs that are semantically closest to the extracted current intent $k_{cur}$. We use an embedding model (e.g., text-embedding-3-small) to compute dense vector representations for all memory units. The relevance score between $k_{cur}$ and each memory unit is calculated using cosine similarity. The top-$K$ memory units $(k_{i},V_{i})$ with the highest scores are selected to form the candidate memory set, $M_{cand}$. This procedure ensures that we only consider memories from sessions with a shared high-level objective, guaranteeing thematic alignment for subsequent stages.
Missing-Slot Guided Filtering.
While the retrieved memories in $M_{cand}$ are thematically aligned with the user’s intent, their immediate utility for the current dialogue turn can vary significantly. To address this, this stage performs a fine-grained filtering, prioritizing memory units that are most likely to resolve the immediate information needs of the dialogue.
Information Gap Identification.
To guide the filtering process, we first must precisely identify what information is required to advance the current task. We leverage an LLM configured as a CoT reasoner (Wei et al. 2022) to analyze the dialogue context $c$ and the overall goal corresponding with intent key $k_{cur}$, so as to enumerate a list of essential task slots that have not yet been filled or confirmed. The CoT prompt is structured to guide the LLM through a logical sequence. First, it enumerates all required slots for the intent, then checks the dialogue history to determine which of these slots have already been filled or confirmed, and finally outputs only those slots that remain unresolved. The result is a list of hypothesized missing slots, $L_{miss}=\{slot_{1},slot_{2},\dots\}$. For example, in a flight booking task where the user has only specified a destination, $L_{miss}$ might be identified as $\{\text{departure\ date},\text{return\ date},\text{seat\ preference}\}$.
Re-ranking by Marginal Slot-Completion Gain.
Having identified the information gaps represented by $L_{miss}$, we then re-rank each QA pair $(q_{i,j},a_{i,j})$ within the candidate set $M_{cand}$ based on its potential to fill one of these gaps. This process is driven by the principle of selecting information with the highest marginal slot-completion gain.
To operationalize this, we fine-tune a smaller, efficient LLaMA-8B model (Meta AI 2024) to act as a specialized filter. This filter estimates the probability that a given QA pair provides an answer for one of the missing slots. For each QA pair $(q_{i,j},a_{i,j})$ from $M_{cand}$, the model computes:
$s_{i,j}=P(y=1\mid c,L_{\text{miss}},q_{i,j},a_{i,j})$ (1)
where $y=1$ signifies that the answer $a_{i,j}$ successfully fills a slot presented in $L_{miss}$. To supervise this filter, we construct a training dataset using the same pipeline as MS-TOD memory bank generation. Specifically, for each held-out session, we simulate missing-slot contexts and label each QA pair from prior sessions as positive (if it fills a missing slot) or negative (otherwise). Detailed training configurations and dataset are provided in Appendix. The model is optimized using a standard binary cross-entropy loss function:
$\mathcal{L}=-\sum_{i,j}\left[y_{i,j}\log s_{i,j}+(1-y_{i,j})\log(1-s_{i,j})\right]$ (2)
To balance the initial semantic relevance from the initial semantic retrieval (denoted as $s^{\text{pre}}_{i,j}$) with the slot-filling utility score $s_{i,j}$ from our filter, we compute a final score:
$s_{\text{final},ij}=\alpha\cdot s^{\text{pre}}_{i,j}+(1-\alpha)\cdot s_{i,j}$ (3)
@@
where $\alpha$ is a hyperparameter. We select the top-$K$ (e,g,. K=5) QA pairs with the highest $s_{\text{final},ij}$ scores.
Response Generation
The final response is generated using the top-$K$ memorys $A_{\text{core}}=\{a_{1},a_{2},\dots,a_{K}\}$, omitting auxiliary questions $q_{i,j}$. An LLM reader receives the dialogue context $c$, the core facts $A_{\text{core}}$, and missing slots $L_{\text{miss}}$ as prompt inputs.
$r=\text{LLM}{\text{Reader}}(\text{prompt}(c,A_{core},L_{\text{miss}}))$ (4)
The prompt instructs the model to: (1) continue the conversation naturally, and (2) proactively address $L_{\text{miss}}$ using $A_{\text{core}}$, e.g., by confirming or suggesting stored values. For example, if memory provided a preferred airline, the system might respond, ”I see you’re flying to Montreal. Last time you flew with Air Canada. Would you like to book with them again?” This proactive strategy, directly informed by our two-stage memory selection, minimizes redundant questions and accelerates task completion. All prompt templates used in our method are provided in Appendix.
Experiments
To enable fine-grained assessment of long-term memory utilization and task completion, we evaluate MemGuide on sessions from the MS-TOD benchmark that are annotated with confirmation-type responses, in which the final utterance explicitly confirms that the user goal has been achieved. Each session is associated with a gold-standard task goal and corresponding slot-value set, enabling precise evaluation using standard metrics of task success and response quality.
Experimental Setups
Evaluation Metrics.